OpenStudy (anonymous):

Simplify (4x/9)+((x-10)/9)

OpenStudy (anonymous):

\[\frac {4x}{9}+\frac {x-10}{9}\] \[\frac {4x+x-10}{9}\] \[\frac {5x-10}{9}\]

OpenStudy (anonymous):

Thank you!!